OpenStudy (anonymous):

Simplify the following expression. 2x − 2y − 5z + 2x + 2y + 3z

OpenStudy (anonymous):

I believe it is x-4y+8z

OpenStudy (anonymous):

these are the answers: 4x − 4y − 2z 4x − 4y − 8z 4x − 4y + 8z 4x − 2z

OpenStudy (anonymous):

okay that helps me

OpenStudy (anonymous):

4x-4y+8z (i think)

OpenStudy (anonymous):

but just make sure with someone else... just to be sure. always double check your work

OpenStudy (anonymous):

is simple 2x+2x=4 -2y+2y=0 -5z+3z=-2z

OpenStudy (anonymous):

oh yeah the -2y and 2y rule each other out. Forgot.

OpenStudy (anonymous):

so its 4x-2z

OpenStudy (anonymous):

you must know the difference of sign

OpenStudy (anonymous):

okay thanks guys

OpenStudy (anonymous):

i didnt see it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

no problem